You are given the following information about a closed economy with no government: Consumption = 115 + 0.6Y Investment = 550 Use the above information to answer the questions that follow:
Q.4.3 Calculate the equilibrium level of income
1
Expert's answer
2020-09-21T08:57:14-0400

At Equilibrium level of Income ;AS=AD.


Y is the income of the economy. Y=C+I


"Y=115+0.6Y+550"


"Y-0.6Y=115+550"


"0.4Y=665"


Y =1,662.5